OpenStudy (anonymous):

I'm doing a geo problem and I've come upon this: JL is 6√6 and JK is 24, but JK is the hypotenuse and 6√6 if I'm not mistaken is 216 because 36*6 is 216...is this just a mistake the teacher made or am I missing something?

OpenStudy (anonymous):

i'll give a medal to the first person who can answer it correctly!

OpenStudy (anonymous):

6*sqrt(6) = 14.7

OpenStudy (anonymous):

no no no. 6 radical 6

OpenStudy (raden):

do you have a diagram ?

OpenStudy (anonymous):

that would be 36 * 6 right?

OpenStudy (anonymous):

yeah hold on

OpenStudy (anonymous):



OpenStudy (anonymous):

36*6 = 216 6*sqrt(6) = 14.7 i'm not quite sure what you are trying to figure out more...?

OpenStudy (anonymous):

look at the diagram

OpenStudy (anonymous):

what are you solving for, the remaingit side?

OpenStudy (anonymous):

oh and i forgot to add the larger triangle is also right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

are there any interiror angle values aside from 90?

OpenStudy (anonymous):

hold on, i just want to know, 6 radical 6 is the same as 36*6

OpenStudy (anonymous):

nope, not the same 36*6 = 216 6*sqrt(6) = 14.7

OpenStudy (anonymous):

but if simplifying the √ 216, wouldn't it turn out to be 6√ 6?

OpenStudy (raden):

yeah, you are right. the simplify of √ 216 is 6√ 6 in other words, if you squres 6√ 6, get 216

OpenStudy (anonymous):

thank you that's what i needed to confirm

OpenStudy (anonymous):

36*6 = 216 6*sqrt(6) = 14.7 (6*sqrt(6))^2 = 216

OpenStudy (raden):

btw, where the position of point J,K,L ? and what are you solving for ?

OpenStudy (anonymous):



OpenStudy (anonymous):

i'm finding JM LM and LK

OpenStudy (anonymous):

i know i have to use pythag theorem, but i just thought it was confusing that the hypotenuse is 24 and the leg is 6√ 6

OpenStudy (raden):

oh, okay. good idea by using the pythagorean theoremt easier to find LK. we can use the formula : LK^2 = JK^2 - JL^2 so, LK^2 = (24)^2 - (6√ 6)^2 simplify the RHS, then solve for LK ...

OpenStudy (anonymous):

thank you

OpenStudy (raden):

and actually, to get JM. we can use the similarities of both triangle are JLM and JLK, see the corresponding of side of them then take the proportion : JL/JK = JM/JL get (6√ 6)/24 = JM/(6√ 6) solve for JM ....

OpenStudy (raden):

after you get JM, then see the triangle of JLM (right triangle). again use the pythagorean theorem : LM^2 = JL^2 - JM^2